# ASH: Animatable Gaussian Splats for Efficient and Photoreal Human Rendering

### v1

@@ 1 Introduction
In more detail, our animatable human avatar is parameterized using Gaussian splats. However, naively learning a mapping from skeletal pose to Gaussian parameters in 3D is intractable given the limited compute budget when constraining ourselves to real-time performance. Thus, we propose to attach the Gaussians onto a deformable mesh template of the human. Given the mesh’s uv parameterization, it allows learning the Gaussian parameters efficiently in 2D texture space. Here, each texel covered by a triangle represents a Gaussian. Thus, the number of Gaussians remains constant, which is in stark contrast to the original formulation. Similarly, we encode the skeletal motion as pose-dependent normal maps. As a result, learning the mapping from skeletal motion to dynamic and controllable Gaussian parameters simplifies to a 2D-to-2D image translation task, which can be efficiently implemented using 2D convolutional architectures. For supervision, we transform the Gaussians into global space using the deformable template and learned Gaussian displacements, splat the Gaussians following the original formulation, and supervise solely on multi-view videos. Our contributions are:
@@ 2 Related Work
In the last few years, volumetric representations [15, 50, 51] and neural radiance fields (NeRF) [35, 74] have received significant attention due to their ability to generate high-quality geometry and appearance [57, 70]. However, rendering a NeRF is typically slow as it requires querying an MLP for each ray sample during volume rendering. To address this, subsequent research focused on accelerating the inference process of NeRF: Neural Sparse Voxel Fields [29] adopts an octree to prune the ray samples. DVGO [56] models the scene with an explicit density and feature grid. Plenoxels [10] and PlenOctree [73] replace the MLP with a hierarchical 3D grid storing spherical harmonics, achieving an interactive test-time framerate. TensoRF [6] and Instant-NGP [36] accelerated inference with compact scene representations, i.e., decomposed tensors and neural hash grids. 3D Gaussian Splatting [22] encodes the scene with Gaussian splats storing the density and spherical harmonics, which achieves state-of-the-art rendering quality and shows real-time capability. However, all the above methods are tailored for static scenes, and it is non-trivial to extend them for modeling the dynamic appearances of clothed humans.
@@
Since this work focuses on animatable human rendering, i.e., at test time, the approach solely takes the skeletal motion as input, we do not discuss works on replay [43, 31, 64, 24, 66, 17], reconstruction [68, 1, 2, 33, 47, 59, 34, 28], and image-based free-viewpoint rendering [62, 45]. Here, according to the underlying shape representation and rendering scheme, we can categorize the literature into two streams, i.e., mesh-based methods and hybrid methods.
Mesh-based methods [71, 5, 61, 49, 13, 69, 3] adopt an explicit, motion-controllable template mesh to model the geometry of clothed humans, with texture space for encoding appearance features. Xu et al. [71] first achieved novel motion and pose synthesis by querying and wrapping texture patches from the captured dataset. Casas et al. [5] and Volino et al. [61] proposed an interactive system that model the appearance as a temporally consistent layered representation in textures space. However, the rendering quality is limited due to the coarse geometry proxy. TNA [49] adopts a texture stack for modeling the dynamic humans’ appearances, though it cannot generate motion-dependent appearances. To address this issue, DDC [13] employs differentiable rendering to learn the non-rigid deformations and dynamic texture maps of clothed humans. At test time, DDC can generalize to novel poses and views and produce real-time photorealistic rendering. Our method outperforms DDC in terms of rendering quality by a large margin while maintaining real-time capability.
Although mesh-based methods provide intuitive control through skeletal poses and integrate seamlessly with established rasterization pipelines, their rendering quality is restrained by the resolution of the template mesh. To this end, hybrid methods are introduced, which articulate the implicit fields with the explicit shape proxy, i.e., parametric human body models [32, 38, 41, 20], or person-specific template meshes. A popular line of research [53, 65, 7, 37, 63, 4, 19, 26, 54, 9, 16] introduced deformable human NeRFs that unwrap the posed space to a shared canonicalized space with inverse kinematics. To better model the pose-dependent appearance of humans, recent studies  [30, 42, 72, 11, 76, 14, 25] further introduce motion-aware residual deformations in the canonicalized space. Neural Actor [30] and HDHumans [14] are most closely related to our work within this category. Neural Actor utilizes the texture map of the parametric human body mesh as local pose features to infer dynamic appearances. However, it fails to generalize to characters with loose outfits. HDHumans jointly optimizes the neural implicit fields and the explicit template mesh and, thus, is able to handle loose clothing. However, both methods are slow and take roughly 5 seconds to render a single frame. In stark contrast, our proposed method is capable of real-time rendering with a quality that is on par with, or even superior to HDHumans.
@@ 3.1 Problem Setting and Background
We assume a segmented multi-view video $\mathbf{I}_{f,c}$ of an actor, recorded in a studio with $C$ synchronized and calibrated cameras, where $f$ and $c$ denote the frame and camera IDs, respectively. $\mathbf{C}_{c}$ denotes the camera projection matrix. Additionally, each frame $\mathbf{I}_{f,c}$ is annotated with the 3D skeletal pose $\boldsymbol{\theta}_{f}\in\mathbb{R}^{D}$ using a markerless motion capture system [58]. Here, $D$ indicates the number of degrees of freedom (DoFs) for the character’s skeleton. The skeletal motion of the subject $\boldsymbol{\bar{\theta}}_{f}\in\mathbb{R}^{k\times D}$ is depicted by a sliding window of skeletal poses from frame $f-k+1$ to frame $f$, where the root translation is normalized w.r.t. the $f$th frame.
For training, our model takes the skeletal poses $\boldsymbol{\theta}_{f}$ and camera parameters $\mathbf{C}_{c}$ as input, renders the animatible Gaussians into image space, and is supervised solely on the multi-view video $\mathbf{I}_{f,c}$. During inference, ASH takes arbitrary skeletal poses $\boldsymbol{\theta}^{\prime}$ and virtual cameras $\mathbf{C}^{\prime}$ as input and generates photorealistic rendering of the subjects at real-time frame rates.
@@
where $\mathcal{N}_{p}$ denotes the set of Gaussian splats covering pixel $\mathbf{p}$. $\alpha^{\prime}_{j}$ refers to the opacity for the $j$th ordered Gaussian splat with respect to the current pixel, i.e. $\alpha^{\prime}_{j}=\alpha_{j}G_{j}(\mathbf{p})$. $H(\cdot)$ indicates the function that converts the spherical harmonics coefficients $\boldsymbol{\eta}_{i}$ and the view direction $\mathbf{d}_{p}$ to an RGB color.
@@ 3.2 Animatable Gaussian Splats
Intuitively, we want to learn a function $\mathcal{F}(\boldsymbol{\bar{\theta}}_{f})=\{\mathcal{G}_{i}\}_{N_{g}}$ that maps the skeletal motion to animatable 3D Gaussian parameters. However, more than $20{,}000$ Gaussian splats are typically required to achieve high-fidelity rendering of clothed humans. Thus, modeling and learning such a function can be challenging, especially when modeling it in 3D. Instead, our idea is to attach the Gaussian splats onto an animatable template mesh of the human, and parameterize the Gaussian splats in 2D texture space, i.e., each texel of the template mesh (covered by a face) stores the parameters of a 3D Gaussian. This enables ASH to efficiently learn the Gaussian parameters in 2D texture space, which we will now describe in more detail.
Animatable Template. To achieve this, we require an animatable human template denoted as $M(\boldsymbol{\theta}_{f})=\mathbf{V}_{f}$, which takes the skeletal motion and computes posed and deformed 3D vertices $\mathbf{V}_{f}$ of a person-specific template mesh $\mathbf{V}_{\mathrm{m}}$. In practice, we leverage the character model of Habermann et al. [13] and refer to the supplemental for further details. To generate the animatable template mesh $\mathbf{V}_{f}$, we first non-rigidly deform the template mesh vertices $\mathbf{V}_{\mathrm{m}}$ in the unposed-canonical space, denoted as $\bar{\mathbf{V}}_{f}$, with skeletal motion-dependent, i.e. $\bar{\boldsymbol{\theta}}_{f}$, and learned embedded derformations [55] and per-vertex displacements. Given the skeletal pose ${\boldsymbol{\theta}}_{f}$, the canonically deformed template mesh vertices $\bar{\mathbf{V}}_{f}$ can then be posed using Dual Quaternion skinning [21], denoted as $\mathbf{V}_{f}$.
@@
Gaussian Texture Decoder. Due to the texel-based 2D parameterization of the 3D Gaussian splats, we can leverage the well-established, efficient 2D convolutional architectures. To formulate the mapping between the 3D skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ and the dynamic Gaussian splats $\{\mathcal{G}_{i}\}_{N}$ on 2D texture space as a image-2-image translation problem [18], we adopt the motion-aware textures $(\mathbf{T}_{\mathrm{n},f},\mathbf{T}_{\mathrm{p},f})$ to depict the 3D skeletal motions $\boldsymbol{\bar{\theta}}_{f}$ in the 2D texture space. The normal textures $\mathbf{T}_{\mathrm{n},f}$ and position textures $\mathbf{T}_{\mathrm{p},f}$ can be computed from the posed deformed template mesh $\mathbf{V}_{f}$ vertices of previous frames through inverse texture mapping. Consequently, we propose motion-aware 2D convolutional neural networks, i.e., the geometry network $\mathcal{E}_{\mathrm{geo}}$, and the appearance network $\mathcal{E}_{\mathrm{app}}$, predicting the geometry and appearance parameters of the Gaussian splats from the motion-aware textures $(\mathbf{T}_{\mathrm{n},f},\mathbf{T}_{\mathrm{p},f})$. The geometry network $\mathcal{E}_{\mathrm{geo}}$ predicts the shape-related parameters, namely, the canonical offset $\mathbf{\bar{d}}_{\mathrm{uv},i}$, scale $\mathbf{s}_{\mathrm{uv},i}$, rotation quaternions $\mathbf{q}_{\mathrm{uv},i}$, and opacity $\mathbf{\alpha}_{\mathrm{uv},i}$:
@@ 3.3 Training Strategy
where $\mathcal{L}_{\mathrm{ssim}}$ denotes the structural similarity index loss measuring the structural difference between two images. $\lambda_{\mathrm{pix}}$ and $\lambda_{\mathrm{ssim}}$ are set to $0.1$ and $0.9$, respectively.
@@ 4.2 Comparisons
Quantitative Comparison. Tab. 1 and Tab. 2 illustrate the quantitative comparison against the competing methods on novel view and pose synthesis. Compared with the real-time capable methods, our method significantly outperforms DDC [13] in PSNR and LPIPS regarding novel-view synthesis, highlighting our method’s superiority in capturing the motion-aware appearances from the training data. In novel pose synthesis, compared to DDC [13], our method demonstrates significantly improved performance. This underscores our method’s generalization ability to novel motions. As for the comparison against the non-real-time approaches, our method consistently surpasses previous works regarding PSNR and LPIPS. Notably, our method is capable of real-time rendering and achieves much better quantitative performances than HDHumans in novel-view synthesis, and comparable performances in novel-pose synthesis.
@@
Qualitative Comparison. Fig. 4 comprises the qualitative comparison on the novel-view and novel-pose rendering: TAVA [26] struggles to handle various motions in the DynaCap dataset [13], resulting in blurry rendering. While NA [30] effectively captures details in subjects wearing tight apparel, it struggles with significant artifacts for subjects in loose outfits. This issue arises from the inherent challenge of representing loose clothing as residual displacements on the parametric human body model. HDHumans [14] stands out among non-real-time competing methods, producing high-fidelity renderings with sharp details. Notably, it preserves the wrinkles and delicate textures on the skirt. However, due to the extensive sampling needed for volume rendering, it takes seconds for HDHumans to render a single frame. In contrast, ASH excels by delivering rendering quality that matches or exceeds HDHumans in real time.
@@ 4.3 Ablations
To assess the effectiveness of the major components of our model, we conduct the following ablative experiments on the novel view and motion synthesis tasks.
@@ 5 Conclusion
In this paper, we introduce ASH, a real-time method for high-quality rendering of animated humans, learned solely from multi-view videos. ASH attaches the 3D Gaussians splats, initially designed for modeling static scenes, onto a deformable mesh template. Bridged by the mesh’s UV parameterization, we can efficiently learn the 3D Gaussians in 2D texture space as a texture translation task. ASH demonstrates significantly better performances quantitatively and qualitatively than state-of-the-art, real-time capable methods on animatable human rendering, and even better performance than the state-of-the-art offline methods. Currently, ASH does not update the underlying deformable template mesh. In the future, we will explore whether the Gaussian splatting can directly improve the 3D mesh geometry.

### v2

@@ 1 Introduction
In more detail, our animatable human avatar is parameterized using Gaussian splats. However, naively learning a mapping from skeletal pose to Gaussian parameters in 3D leads to inferior quality when constraining ourselves to real-time performance. Thus, we propose to attach the Gaussians onto a deformable mesh template of the human. Given the mesh’s uv parameterization, it allows learning the Gaussian parameters efficiently in 2D texture space. Here, each texel covered by a triangle represents a Gaussian. Thus, the number of Gaussians remains constant, which is in stark contrast to the original formulation. Similarly, we encode the skeletal motion as pose-dependent normal maps. As a result, learning the mapping from skeletal motion to dynamic and controllable Gaussian parameters simplifies to a 2D-to-2D image translation task, which can be efficiently implemented using 2D convolutional architectures. For supervision, we transform the Gaussians into global 3D space using the deformable template and learned Gaussian displacements, splat the Gaussians following the original formulation, and supervise solely on multi-view videos. Our contributions are:
@@ 2 Related Work
In the last few years, volumetric representations [18, 58, 59] and neural radiance fields (NeRF) [42, 83] have received significant attention due to their ability to generate high-quality geometry and appearance [65, 79]. However, rendering a NeRF is typically slow as it requires querying an MLP for each ray sample during volume rendering. To address this, subsequent research focused on accelerating the inference process of NeRF: Neural Sparse Voxel Fields [34] adopts an octree to prune the ray samples. DVGO [64] models the scene with an explicit density and feature grid. Plenoxels [10] and PlenOctree [82] replace the MLP with a hierarchical 3D grid storing spherical harmonics, achieving an interactive test-time framerate. TensoRF [6] and Instant-NGP [43] achieved faster inference with compact scene representations, i.e., decomposed tensors and neural hash grids. 3D Gaussian Splatting [26] encodes the scene with Gaussian splats storing the density and spherical harmonics, which achieves state-of-the-art rendering quality and shows real-time capability. However, all the above methods are tailored for static scenes, and it is non-trivial to extend them for modeling the dynamic appearances of clothed humans.
@@
Since this work focuses on animatable human rendering, i.e., at test time, the approach solely takes the skeletal motion as input, we do not discuss works on replay [50, 36, 73, 28, 75, 20], reconstruction [77, 1, 2, 39, 54, 67, 41, 33, 30, 23, 16, 14, 13], and image-based free-viewpoint rendering [70, 52, 56]. Here, according to the underlying shape representation and rendering scheme, we can categorize the literature into two streams, i.e., mesh-based methods and hybrid methods.
Mesh-based methods [80, 5, 69, 57, 15, 78, 3] adopt an explicit, motion-controllable template mesh to model the geometry of clothed humans, with texture space for encoding appearance features. Xu et al. [80] first achieved novel motion and pose synthesis by querying and wrapping texture patches from the captured dataset. Casas et al. [5] and Volino et al. [69] proposed an interactive system that models the appearance as a temporally consistent layered representation in textures space. However, the rendering quality is limited due to the coarse geometry proxy. TNA [57] adopts a texture stack for modeling the dynamic humans’ appearances, though it cannot generate motion-dependent appearance. To address this issue, DDC [15] employs differentiable rendering to learn the non-rigid deformations and dynamic texture maps of clothed humans. At test time, DDC can generalize to novel poses and views and produce real-time photorealistic renderings. Our method outperforms DDC in terms of rendering quality by a large margin while maintaining real-time capability.
Although mesh-based methods provide intuitive control through skeletal poses and integrate seamlessly with established rasterization pipelines, their rendering quality is restrained by the resolution of the template mesh. To this end, hybrid methods are introduced, which articulate the implicit fields with the explicit shape proxy, i.e., parametric human body models [37, 45, 48, 24], or person-specific template meshes. A popular line of research [61, 74, 7, 44, 71, 4, 22, 31, 62, 9, 19] introduced deformable human NeRFs that unwrap the posed space to a shared canonicalized space with inverse kinematics. To better model the pose-dependent appearance of humans, recent studies  [35, 49, 81, 11, 85, 17, 29, 86] further introduce motion-aware residual deformations in the canonicalized space. Neural Actor [35] and HDHumans [17] are most closely related to our work within this category. Neural Actor utilizes the texture map of the parametric human body mesh as local pose features to infer dynamic appearances. However, it fails to generalize to characters with loose outfits. HDHumans jointly optimizes the neural implicit fields and the explicit template mesh and, thus, is able to handle loose clothing. However, both methods are slow and take roughly 5 seconds to render a single frame. In stark contrast, our proposed method is capable of real-time rendering with a quality on par with or even superior to HDHumans.
@@ 3.1 Problem Setting and Background
We assume a segmented multi-view video $\mathbf{I}_{f,c}$ of an actor, recorded in a studio with $C$ synchronized and calibrated cameras, where $f$ and $c$ denote the frame and camera IDs, respectively. $\mathbf{C}_{c}$ denotes the camera projection matrix. Additionally, each frame $\mathbf{I}_{f,c}$ is annotated with the 3D skeletal pose $\boldsymbol{\theta}_{f}\in\mathbb{R}^{D}$ using a markerless motion capture system [66]. Here, $D$ indicates the number of degrees of freedom (DoFs) for the character’s skeleton. The skeletal motion of the subject $\boldsymbol{\bar{\theta}}_{f}\in\mathbb{R}^{k\times D}$ is depicted by a sliding window of skeletal poses from frame $f-k+1$ to frame $f$ where the root translation is normalized w.r.t. the $f$th frame.
For training, our model takes the skeletal poses $\boldsymbol{\theta}_{f}$ and camera parameters $\mathbf{C}_{c}$ as input, renders the animatible Gaussians into image space, and is supervised solely on the multi-view video $\mathbf{I}_{f,c}$. During inference, ASH takes arbitrary skeletal poses $\boldsymbol{\theta}^{\prime}$ and virtual cameras $\mathbf{C}^{\prime}$ as input and generates photorealistic rendering of the subjects at real-time frame rates ($\mathbf{29.64fps}$). The detailed runtime breakdown is reported in the appendix.
@@
where $\mathcal{N}_{p}$ denotes the set of Gaussian splats covering pixel $\mathbf{p}$. $\alpha^{\prime}_{j}$ refers to the opacity for the $j$th ordered Gaussian splat with respect to the current pixel, i.e. $\alpha^{\prime}_{j}=\alpha_{j}G_{j}(\mathbf{p})$. $H(\cdot)$ indicates the function that converts the spherical harmonics coefficients $\boldsymbol{\eta}_{i}$ and the view direction $\mathbf{d}_{\mathbf{p}}$ to an RGB color.
@@ 3.2 Animatable Gaussian Splats
Intuitively, we want to learn a function $\mathcal{F}(\boldsymbol{\bar{\theta}}_{f})=\{\mathcal{G}_{i}\}_{N_{g}}$ that maps the skeletal motion to animatable 3D Gaussian parameters. However, more than $20{,}000$ Gaussian splats are typically required to achieve high-fidelity renderings of clothed humans. Thus, modeling and learning such a function can be challenging, especially when modeling it in 3D. Instead, our idea is to attach the Gaussian splats onto an animatable template mesh of the human, and parameterize the Gaussian splats in 2D texture space, i.e., each texel of the template mesh (covered by a face) stores the parameters of a 3D Gaussian. This enables ASH to efficiently learn the Gaussian parameters in 2D texture space, which we will now describe in more detail.
Animatable Template. To achieve this, we require an animatable human template denoted as $M(\boldsymbol{\theta}_{f})=\mathbf{V}_{f}$, which takes the skeletal motion and computes posed and deformed 3D vertices $\mathbf{V}_{f}$ of a person-specific template mesh $\mathbf{V}_{\mathrm{m}}$. In practice, we leverage the character model of Habermann et al. [15] and refer to the appendix for further details. To generate the animatable template mesh $\mathbf{V}_{f}$, we first non-rigidly deform the original template mesh vertices $\mathbf{V}_{\mathrm{m}}$ in the unposed-canonical space, denoted as $\bar{\mathbf{V}}_{f}$, with skeletal motion-dependent, i.e. $\bar{\boldsymbol{\theta}}_{f}$, and learned embedded derformations [63] and per-vertex displacements. Given the skeletal pose ${\boldsymbol{\theta}}_{f}$, the canonically deformed template mesh vertices $\bar{\mathbf{V}}_{f}$ can then be posed using Dual Quaternion skinning [25], denoted as $\mathbf{V}_{f}$.
@@
Gaussian Texture Decoder. Due to the texel-based 2D parameterization of the 3D Gaussian splats, we can leverage the well-established, efficient 2D convolutional architectures. To formulate the mapping between the 3D skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ and the dynamic Gaussian splats $\{\mathcal{G}_{i}\}_{N}$ on 2D texture space as a image-2-image translation problem [21], we adopt the motion-aware textures $(\mathbf{T}_{\mathrm{n},f},\mathbf{T}_{\mathrm{p},f})$ to depict the 3D skeletal motions $\boldsymbol{\bar{\theta}}_{f}$ in the 2D texture space. The normal textures $\mathbf{T}_{\mathrm{n},f}$ and position textures $\mathbf{T}_{\mathrm{p},f}$ can be computed from the posed and deformed template mesh $\mathbf{V}_{f}$ vertices through inverse texture mapping. Consequently, we propose motion-aware 2D convolutional neural networks, i.e., the geometry network $\mathcal{E}_{\mathrm{geo}}$, and the appearance network $\mathcal{E}_{\mathrm{app}}$, predicting the geometry and appearance parameters of the Gaussian splats from the motion-aware textures $(\mathbf{T}_{\mathrm{n},f},\mathbf{T}_{\mathrm{p},f})$. The geometry network $\mathcal{E}_{\mathrm{geo}}$ predicts the shape-related parameters, namely, the canonical offset $\mathbf{\bar{d}}_{\mathrm{uv},i}$, scale $\mathbf{s}_{\mathrm{uv},i}$, rotation quaternions $\mathbf{q}_{\mathrm{uv},i}$, and opacity $\mathbf{\alpha}_{\mathrm{uv},i}$:
@@ 3.3 Training Strategy
where $\mathcal{L}_{\mathrm{ssim}}$ denotes the structural similarity index loss [72] measuring the structural difference between two images. $\lambda_{\mathrm{pix}}$ and $\lambda_{\mathrm{ssim}}$ are set to $0.1$ and $0.9$, respectively.
@@ 4.2 Comparisons
Quantitative Comparison. Tab. 1 and Tab. 2 illustrate the quantitative comparison against the competing methods on novel view and pose synthesis. Compared with the real-time capable methods, our method significantly outperforms DDC [15] in PSNR and LPIPS regarding novel-view synthesis, highlighting our method’s superiority in capturing the motion-aware appearances from the training data. In novel pose synthesis, compared to DDC [15], our method demonstrates significantly improved performance. This underscores our method’s generalization ability to novel motions. As for the comparison against the non-real-time approaches, our method consistently surpasses previous works regarding PSNR and LPIPS. Notably, our method is capable of real-time rendering and achieves remarkably better quantitative accuracy than HDHumans in novel-view synthesis, and comparable performances in novel-pose synthesis.
@@
Qualitative Comparison. Fig. 4 comprises the qualitative comparison on the novel-view and novel-pose rendering: TAVA [31] struggles to handle various motions in the DynaCap dataset [15], resulting in blurry renderings. While NA [35] effectively captures details for subjects wearing tight apparel, it struggles with significant artifacts for subjects in loose outfits. This issue arises from the inherent challenge of representing loose clothing as residual displacements on the parametric human body model. HDHumans [17] stands out among non-real-time competing methods, producing high-fidelity renderings with sharp details. However, due to the extensive sampling needed for volume rendering, it takes seconds for HDHumans to render a single frame. In contrast, ASH excels by delivering rendering quality that matches or exceeds HDHumans’ quality in real-time.
@@ 4.3 Ablations
To assess the effectiveness of the major components of our method, we conduct the following ablative experiments on the novel view and motion synthesis tasks.
@@ 5 Conclusion
In this paper, we introduce ASH, a real-time method for high-quality rendering of animated humans, learned solely from multi-view videos. ASH attaches the 3D Gaussians splats, initially designed for static scenes, onto a deformable mesh template. Bridged by the mesh’s UV parameterization, we can efficiently learn the 3D Gaussians in 2D texture space as an image-2-image translation task. ASH demonstrates significantly better performances quantitatively and qualitatively than state-of-the-art, real-time capable methods on animatable human rendering, and even better performance than the state-of-the-art offline methods. Currently, ASH does not update the underlying deformable template mesh. In the future, we will explore whether the Gaussian splatting can directly improve the 3D mesh geometry.
6 Acknowledgement
Christian Theobalt was supported by ERC Consolidator Grant 4DReply (No.770784). Adam Kortylewski was supported by the German Science Foundation (No.468670075). This project was also supported by the Saarbrucken Research Center for Visual Computing, Interaction, and AI.